OpenStudy (meilendurcer):

1 3/5 divide 5/8 = A. 1 B. 1 1/5 C. 2 14/25 D. 2 2/5 i have A.

OpenStudy (anonymous):

A.11/5

OpenStudy (meilendurcer):

you mean 1 ,1/5?

hero (hero):

\[1 \frac{3}{5} \div \frac{5}{8} \rightarrow 1\frac{3}{5} \times \frac{8}{5} \]

OpenStudy (meilendurcer):

so i multiple those?

hero (hero):

Hang on

OpenStudy (meilendurcer):

ok

hero (hero):

\[\left(1+\frac{3}{5}\right) \left(\frac{8}{5}\right) = \frac{8}{5} + \frac{24}{25} = \frac{40}{25} + \frac{24}{25} = \frac{64}{25} = 2 \frac{14}{25}\]